Circular No. 7807

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

(26308) 1998 SM_{165}

M. E. Brown and C. A. Trujillo, California Institute of Technology, report the discovery of a companion to the transneptunian object (26308): “Observations were collected on two dates, spanning ∼ 45 min and centered at 2001 Dec. 22.536 and 28.411 UT, as part of an ongoing survey for transneptunian binaries undertaken with the Hubble Space Telescope STIS instrument. Separations of 0′′.2056 ± 0′′.0009 in p.a. 299^{o}.5 ± 0^{o}.2 and 0′′.2334 ± 0′′.0012 in p.a. 267^{o}.2 ± 0^{o}.3 were measured for the two dates, respectively. At a mean geocentric distance of 34.9 AU, this separation corresponds to at least 6000 km between components. The pair exhibit a mean flux difference of ∼ 1.9 magnitudes in the unfiltered images. The flux has only been roughly estimated because the lightcurve of the primary (Romanishin et al. 2001, Bull. AAS 33, 1031) was evident but incomplete in the discovery data.”

1999 TC_{36}

F. Marchis, University of California, Berkeley (UCB); and J. Berthier, Institut de Mécanique Céleste, Centre National de la Recherche Scientifique, report for their collaboration (H. Boenhardt, O. Hainaut, A. Delsanti, European Southern Observatory; I. de Pater, UCB; C. Dumas, Jet Propulsion Laboratory; D. Gavel, Lawrence Livermore National Laboratory) the detection of the companion of the plutino 1999 TC_{36} (cf. IAUC 7787) using the Adaptive Optics on the Lick Shane 3-m telescope: “After careful analysis, observations taken on 2001 Oct. 4.315 during an appulse

event with a Tycho-2 star (cf. http://astron.berkeley.edu/~fmarchis/

Science/TNOs Appulse/), we detect, on K-band images with 0′′.22 angular resolution, the companion at a separation of 0′′.246 ± 0′′.017 at p.a. 322^{o} ± 1^{o}. The difference of magnitude is estimated to be 1.89 ± 0.14. We encourage other groups to observe the next appulse events on 2002 Jan. 30, Feb. 2 and Oct. 10. Astrometric measurement is also needed, since 1999 TC_{36} was ∼ 10′′ from its expected position.”

COMET C/1996 Y2 (SOHO) MPEC 2002-B01 contains astrometry and orbital elements for a comet

visible in C2 and C3 and found by X. Leprette on SOHO website images:

1996 UT  α_{2000}  δ_{2000}
Dec. 29.990  18^{h}49.^{m}1  −25^{o}33′
2002 January 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green